OpenStudy (anonymous):

What is the center of the ellipse defined by the equation (x+8)^2/36 + (y-7)^2/9 = 1?

OpenStudy (turingtest):

an ellipse of the form\[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\]has center\[(h,k)\]

OpenStudy (anonymous):

A. (-8, -7) B. (-8, 7) C. (8, -7) D. (8, 7)

OpenStudy (turingtest):

use the info I gave you and this will be easy